The triangles are similar, find the missing side. 2/3=4/6=5/x, the missing side is x.
Answer by stanbon(75887)

4/6 = 5/x
Coss-multiply to get:
4x = 30
x = 30/4 = 15/2 = 7.5
Cheers,
Stan H.